40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 12 01:09:56 2016, gautam, Update, LSC, Recovering DRMI locking 
    Reply  Wed Nov 16 03:10:01 2016, gautam, Update, LSC, DRMI locked on 1f and 3f signals 
       Reply  Wed Nov 16 08:14:43 2016, Steve, Update, LSC, DRMI locked on 1f and 3f signals 5hrs.png
       Reply  Mon Nov 21 14:02:32 2016, gautam, Update, LSC, DRMI locked on 3f signals, arms held on ALS DRMIArms_Nov20.pdf
    Reply  Wed Nov 23 16:21:02 2016, gautam, Update, LSC, ITMY UL glitches are back ITMYproblemsBack.png
       Reply  Mon Nov 28 10:27:13 2016, gautam, Update, SUS, ITMY UL glitches are back ITMY_satboxSpectra.pdfITMY_UL_testerBox.png
       Reply  Tue Nov 29 11:07:37 2016, Steve, Update, LSC, ITMY UL glitches are back gliching400d.png
          Reply  Wed Nov 30 01:47:56 2016, gautam, Update, LSC, Suspension woes IMCwoes.png
             Reply  Wed Nov 30 17:08:56 2016, gautam, Update, LSC, Binary output breakout box removed 
                Reply  Thu Dec 1 02:19:13 2016, gautam, Update, LSC, Binary output breakout box restored MC1_glitch_fast.pngITMY_testerGlitch.pngITMYsatbox.pngULcomparison.pdf
                   Reply  Thu Dec 1 08:02:57 2016, Steve, Update, LSC, glitching ITMY_UL_LL ITMY_UL_LL.png
                      Reply  Fri Dec 2 11:56:42 2016, gautam, Update, LSC, MC1 LEMO jiggled IMG_3474.JPGMC1_LR_dead.png
                         Reply  Mon Dec 5 15:05:37 2016, gautam, Update, LSC, MC1 glitches are back Mc1glitches.png
                      Reply  Thu Dec 8 10:13:43 2016, Steve, Update, LSC, glitching ITMY_UL has a history glitching__ITMY-UL_2007.png
Message ID: 12630     Entry time: Mon Nov 21 14:02:32 2016     In reply to: 12619
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locked on 3f signals, arms held on ALS 

Over the weekend, I was successful in locking the DRMI with the arms held on ALS. The locks were fairly robust, lasting order of minutes, and was able to reacquire by itself when it lost the lock in <1min. I had to tweak the demod phases and loop gains further compared to the 1f lock with no arms, but eventually I was able to run a sensing matrix measurement as well. A summary of the steps I had to follow:

  • Lock on 1f signals, no arms, and run sensing lines, adjust REFL33 and REFL 165 demod phases to align PRCL, MICH and SRCL as best as possible to REFL33I, REFL165Q and REFL165I respectively
  • I also set the offsets to the 'B' inputs at this stage
  • Lock arms on ALS, engage DRMI locking on 3f signals (the restore script resets some values like the 'B' channel offsets, so I modified the restore script to set the offsets I most recently measured)
  • I was able to achieve short locks on the settings from the locking with no arms - I set the loop gains using the UGF servos and ran some sensing lines to get an idea of what the final demod phases should be
  • Adjusted the demod phases, locked the DRMI again (with CARM offset = -4.0), and took another sensing matrix measurement (~2mins). The data was analyzed using the set of scripts EricQ has made for this purpose, here is the result from a lock yesterday evening (the radial axis is meant to be demod board output volts per meter but the calibration I used may be wrong)

I've updated the appropriate fields in the restore script. Now that the DRMI locking is somewhat stable again, I think the next step towards the full lock would be to zero the CARM offset and turning on the AO path.

On the downside, I noticed yesterday that ITMY UL shadow sensor readback was glitching again - for the locking yesterday, I simply held the output of that channel to the input matrix, which worked fine. I had already done some debugging on the Sat. Box with the help of the tester box, but unlike the PRM sat. box, I did not find anything obviously wrong with the ITMY one... I also ran into a CDS issue when I tried to run the script that sets the phase tracker UGF - the script reported that the channels it was supposed to read (the I and Q outputs of the ALS signal, e.g. C1:ALS-BEATX_FINE_I_OUT) did not exist. The same channels worked on dataviever though, so I am not sure what the problem was. Some time later, the script worked fine too. Something to look out for in the future I guess..

Attachment 1: DRMIArms_Nov20.pdf  71 kB  | Hide | Hide all
DRMIArms_Nov20.pdf
ELOG V3.1.3-